A body moving with a uniform acceleration has a velocity of 12.0cm/s when its x coordinate is 3.00cm if it's coordinate 2.00s later is -5.00cm, what is the magnitude of it's acceleration.
1
Expert's answer
2019-09-13T10:16:44-0400

Task #Q94392

"v=12m\/s"

"x_1=3sm"

"x_2=-5sm"

"t=2s"

______________

"a-?"

______________


"S=x_1-x_2=3-(-5)=8"

"S=vt+(at^2)\/2"

"8=12*2+(a*4)\/2"

"a=-8m\/s^2"